# Ecosystem-level Analysis of Deployed Machine Learning Reveals Homogeneous Outcomes

### v1

@@ 4 Do Model Improvements Improve Systemic Failures?
Ecosystem-level analysis provides a new means for understanding how models change and how those changes will impact people. When these models change, what are the broader consequences? Do single-model improvements on average improve ecosystem-level outcomes by reducing systemic failures? To what extent are the individuals for whom the model improves those that were previously systemically failed?
@@
Chen et al. (2022a) evaluated the performance of the commercial APIs on the same 11 evaluation datasets each year in 2020–2022. Of all year-over-year comparisons, we restrict our attention to cases where one of the three APIs for a given task improves by at least 0.5% accuracy. [^4] The supplement contains an analysis that confirms our findings are robust to this threshold. Let $h_{\text{imp}}$ denote the model that improved. We identify the instances that $h_{\text{imp}}$ initially misclassified in the first year as potential improvements and the subset of these instances that $h_{\text{imp}}$ correctly classified in the second year as improvements. Considering the initial distribution of failures for $h_{\text{imp}}$, we can ask where does the $h_{\text{imp}}$ improve? We answer this by comparing the distribution of outcome profiles for the other models (besides $h_{\text{imp}}$) between the potential improvement and improvement sets.
@@
Figure 4: Model improvement is not concentrated on systemic failures. When a model improves, we compare the distribution of outcome profiles of the other two models on its initial failures (potential improvements) to the distribution on the instances it improved on (observed improvements). Across all improvements, including Amazon’s improvement on waimai (left), there is a clear over-improvement on [✓, ✓] (above $y=x$ on right) and under-improvement on [X, X] (below $y=x$ on right).
@@
As a case study, we consider Amazon’s improvement on the waimai dataset from 2020 to 2021. In 3(a), from left to right, [X, X] indicates the other APIs (Baidu, Google) both fail, [X, ✓] and [✓, X] indicate exactly one of the other APIs succeed, and [✓, ✓] indicates both succeed. The majority (64.7%) of Amazon’s improvement is on instances already classified correctly by the other two APIs, far exceeding the fraction of potential improvements that were classified correctly by Baidu and Google (34.1%) in 2020. In contrast, for systemic failures, the improvement of 11.6% falls far short of the potential improvement of 36.7%. In fact, since models can also fail on instances they previously classified correctly, the model’s improvement on systemic failures is even worse in terms of net improvement. Amazon’s improvement amounts to no net reduction of systemic failures: the model improves on 78 systemic failures but also regresses on 78 instances that become new systemic failures, amounting to no net improvement.
This finding is not unique to Amazon’s improvement on the waimai dataset: across all periods of model improvement that we observe from 2020-2022 on the 11 datasets, we observe the same pattern of homogeneous improvements. In 3(b), we compare the observed improvement distribution [^5] The supplement contains analysis comparing ‘net improvements’ to ‘potential improvements’ as well; the trends are consistent across both analyses. ($y$ axis) to the potential improvement distributions ($x$ axis) across all model improvements. We find a clear pattern: systemic failures (the [X, X] category) consistently represent far less of the observed improvement set compared to the potential improvement set, indicating that when models improve, they under-improve on users that are already being failed by other models. Instead, model improvements especially concentrate on instances where both other models succeeded already.
Ecosystem-level analysis in the context of model improvements disambiguates two plausible situations that are otherwise conflated: does single-model improvement (i) marginally or (ii) substantively reduce systemic failures? We find the reduction of systemic failures is consistently marginal: in every case, the reduction fails to match the the distribution seen in the previous year (i.e. every [X, X] red point is below the line in 3(b)).
@@ 5 Ecosystem-level Analysis in Dermatology (DDI)
Having demonstrated how ecosystem-level analysis reveals homogeneous outcomes across machine learning deployments, we apply the ecosystem-level methodology to medical imaging. Machine learning mediates consequential decisions in medical imaging: for example, models make predictions that inform high-stakes treatment decisions by dermatologists.
@@ 5.1 Data
Daneshjou et al. (2022) introduced the Diverse Dermatology Images (DDI) dataset of 656 skin lesion images to evaluate binary classification performance at detecting whether a lesion was malignant or benign. Images were labelled with the ground-truth based on confirmation from an external medical procedure (biopsy). In addition, each image is annotated with skintone metadata using the Fitzpatrick scale according to one of three categories: Fitzpatrick I & II (light skin), Fitzpatrick III & IV (medium skin), and Fitzpatrick V & VI (dark skin). We use the predictions from Daneshjou et al. (2022) on DDI for two prominent ML models (ModelDerm (Han et al., 2020) and DeepDerm (Esteva et al., 2017)) and two board-certified dermatologists. [^6] Daneshjou et al. (2022) also evaluated a third model (HAM10K; Tschandl et al., 2018) that almost always predicts the majority class in this class-imbalanced setting. We exclude this model since its failures are not interesting, but replicate our analyses in the supplement to show the findings still hold if it is included. Further details are deferred to the supplement.
@@ 5.2 Results
Figure 5: Profile polarization for models and humans. Consistent with HAPI, model predictions (left) yield polarized profiles on DDI. Human predictions (right) are even more polarized than models.
Finding 3: Both humans and models yield polarized profiles; humans are more polarized.
We compare observed and expected ecosystem-level outcomes on DDI for models (4(a)) and humans (4(b)). Consistent with the general trends in HAPI, model predictions yield polarized profiles. For human predictions from board-certified dermatologists, we also see that profiles are polarized. However, in comparing the two, we find that humans yield even more polarized profiles. We take this to be an important reminder: while we may intuit models in practice are likely to produce homogeneous outcomes for a range of reasons, we should expect humans to also produce homogeneous outcomes that, at times, might be even more homogeneous.
@@
Figure 6: Racial disparities for models but not humans. We stratify ecosystem-level analysis in DDI by the three skin tone categories, plotting the difference between observed and expected rates. Models (left) show clear racial disparities, exhibiting the most polarization for the darkest skin tones, whereas humans (right) show no significant racial disparity.
@@
Standard analyses of machine learning’s impact focus on model performance across groups (e.g. Buolamwini and Gebru, 2018; Koenecke et al., 2020). In AI for medicine, several works have taken such an approach towards understanding fairness (Obermeyer et al., 2019; Seyyed-Kalantari et al., 2021; Kim et al., 2022; Colwell et al., 2022). Daneshjou et al. (2022) demonstrate racial disparities for both model and human predictions on DDI, and predictive performance is worst for darkest skin tones, aligning with broader trends of racial discrimination in medicine and healthcare (Vyas et al., 2020; Williams and Wyatt, 2015; Fiscella and Sanders, 2016).
Ecosystem-level analysis can build on these efforts: we conduct the same analysis from Figure 5 stratified by skin tone. In Figure 6, we plot the difference between the observed and expected rates on the $y$ axis: the All bars (blue) reproduce the polarization results from Figure 5. Strikingly, ecosystem-level-analysis surfaces an important contrast between model behavior (left) and human behavior (right). Models (5(a)) are most polarized for darkest skin tones (Fitzpatrick V & VI; dark brown) and slightly anti-polarized for the lightest skin tones (Fitzpatrick I & II; cream). In contrast, humans (5(b)) show no significant variation as a function of skin tone. Ecosystem-level analysis therefore identifies a new form of racial disparity not previously documented. Critically, while prior works show racial disparities for both models and humans (to differing extents), here we find racial disparity that is salient for models but not present for humans.
@@ 6.1 Explanations for Homogeneous Outcomes
We posit properties of the data may give rise to homogeneous outcomes: we explore three such properties.
@@
First, human ambiguity on the ground-truth is one notion of example difficulty that may predict homogeneous outcomes. To test this, we make use of the 10 human annotations per example in the fer+ dataset within HAPI: we find that the systemic failure rate is monotonically correlated with annotator disagreement with the majority label. However, we find that profiles are polarized for all amounts of disagreement, even the most ambiguously labeled images, indicating human ambiguity is only partially explanatory of ecosystem-level model outcomes.
@@
Second, human error (when the ground-truth is externally determined) is another notion of example difficulty that may predict homogeneous outcomes. To test this, we make use of human predictions on DDI: we stratify ecosystem-level analysis by dermatologist performance: (i) instances they both get right, (ii) instances they both get wrong, and (iii) instances where they disagree and exactly one is right. We find that when both dermatologists fail, there continues to be profile polarization, including the outcome of both models succeeding more often than expected. However, when both dermatologists succeed, there is no polarization and the observed rates almost exactly match the expected rates for every outcome profile. This suggests human error is also partially predictive of ecosystem-level model outcomes.
@@
Finally, more expressive theoretical models can better capture the observed trends than our simple full instance-level independence model. We introduce a two-parameter model. $\alpha$ fraction of instances are categorized as difficult and the remaining $1-\alpha$ are easy. A model’s failure rate $\bar{f_{j}}$ over all examples scales to $(1+\Delta)\bar{f_{j}}$ on hard examples and $\left(1-\frac{\alpha\Delta}{1-\alpha}\right)\bar{f_{j}}$ on easy examples. Partitioning examples in this way, while continuing to assume instance-level independence, inherently homogenizes: models are more likely to systemically succeed on easy instances and systemically fail on hard instances. To best fit the HAPI data, the resulting average $\alpha$ ($0.2-0.3$) and $\Delta$ ($1-4$, meaning these examples are 25×2-5$\times$2 - 5 × harder) values are quite extreme. In other words, for this theoretical model to fit the data, a significant fraction ($\approx 25\%$) would need to be considerably harder (3.5$\approx$ 3.5$\times$ 3.5 ×) than the overall performance.
@@
These analyses contribute to an overarching hypothesis that example difficulty partially explains homogeneous outcomes. While we discuss the construct of difficulty in the supplement, we draw attention to two points. First, difficulty is largely in the eye of the beholder: what is perceived as difficult to humans or models can differ (e.g. adversarial examples; Goodfellow et al., 2014) and is ultimately a constructed view of data and the world. Second, whether or not homogeneous outcomes are caused by example difficulty does not change their societal impact. The consequences of systemic rejection can be material and serious (e.g. unemployment).
@@
6.2 Implications
Our work provides new opportunities for researchers and invites new priorities for policymakers.
@@ 6.2 Implications
Our paper shows that ecosystem-level research can reveal previously-invisible social impacts of machine learning. We believe this methodology concretizes the impact of decision-making in real contexts in which individuals would typically be affected by decisions from many actors (e.g. job applications, medical treatment, loan applications, or rent pricing). As we demonstrate in DDI, our methodology applies equally to human-only, machine-only, and more complex intertwined decision-making: we anticipate understanding of homogeneous outcomes arising from any or all of these sources will be valuable.
@@
Given the pervasive and persisting homogeneous outcomes we document, there may be a need for policy intervention. Since in many cases no single decision-maker can observe the decisions made by others in the ecosystem, homogeneous outcomes are not currently the responsibility of any single decision-maker. Consequently, policy could implement mechanisms to better monitor ecosystem-level outcomes and incentivize ecosystem-level improvement. In parallel, regulators should establish mechanisms for recourse or redress for those currently systemically failed by machine learning (see Voigt and von dem Bussche, 2017; Cen and Raghavan, 2023).
@@ 7 Conclusion
We introduce ecosystem-level analysis as a new methodology for understanding the cumulative societal impact of machine learning on individuals. Our analysis on HAPI establishes general trends towards homogeneous outcomes that are largely unaddressed even when models improve. Our analysis on DDI exposes new forms of racial disparity in medical imaging that arise in model predictions but not in human predictions. To ensure machine learning advances the public interest, we should use approaches like ecosystem-level analysis to holistically characterize its impact.
@@ Appendix A Data-centric Explanations for Homogeneous Outcomes
We test three variants of this hypothesis. In § A.1, we examine how the level of annotator disagreement in the ground truth label impacts ecosystem-level behavior. In § A.2, we test how the accuracy of human dermatologists in predicting the malignancy of a skin lesion image correlates with model profile polarization. Finally, to build on these finer-grained empirical analyses, in § A.3, we introduce a more express theoretical model. Under this model, parameterized by two difficulty parameters, we compute a different expected rate for ecosystem-level outcomes, showing it can better recover the observation distribution.

### v2

@@ 4 Do Model Improvements Improve Systemic Failures?
Ecosystem-level analysis provides a new means for understanding how models change and how those changes impact people. When models change, what are the broader consequences across their model ecosystem? Do single-model improvements on average improve ecosystem-level outcomes by reducing systemic failures? And to what extent are the individuals for whom the model improves the same individuals were previously systemically failed?
@@
Chen et al. (2022a) evaluated the performance of the commercial APIs on the same eleven evaluation datasets each year in 2020–2022. Of all year-over-year comparisons, we restrict our attention to cases where one of the three APIs for a given task improves by at least 0.5% accuracy. [^3] The supplement § B.3 contains an analysis that confirms our findings are robust to alternate thresholds. Let $h_{\text{imp}}$ denote the model that improved. We identify the instances that $h_{\text{imp}}$ initially misclassified in the first year as potential improvements and the subset of these instances that $h_{\text{imp}}$ correctly classified in the second year as improvements. Considering the initial distribution of failures for $h_{\text{imp}}$, we can ask where does the $h_{\text{imp}}$ improve? We answer this by comparing the distribution of outcome profiles for the other models (besides $h_{\text{imp}}$) between the potential improvement and improvement sets.
@@
Figure 4: Model improvement is not concentrated on systemic failures. When a model improves, we compare the distribution of outcome profiles of the other two models on its initial failures (potential improvements) to the distribution on the instances it improved on (observed improvements). Across all improvements, including Amazon’s improvement on waimai (left), there is a clear over-improvement on [✓, ✓] (above $y=x$ on right) and under-improvement on [X, X] (below the identity line on right).
@@
As a case study, we consider Amazon’s improvement on the waimai dataset from 2020 to 2021. In 4(a), from left to right, [X, X] indicates the other APIs (Baidu, Google) both fail, [X, ✓] and [✓, X] indicate exactly one of the other APIs succeed, and [✓, ✓] indicates both succeed. The majority (64.7%) of Amazon’s improvement is on instances already classified correctly by the other two APIs, far exceeding the fraction of potential improvements that were classified correctly by Baidu and Google (34.1%) in 2020. In contrast, for systemic failures, the improvement of 11.6% falls far short of the potential improvement of 36.7%. In fact, since models can also fail on instances they previously classified correctly, the model’s improvement on systemic failures is even worse in terms of net improvement. Amazon’s improvement amounts to no net reduction of systemic failures: the model improves on 78 systemic failures but also regresses on 78 instances that become new systemic failures, amounting to no net improvement. The Baidu and Google APIs similarly show little improvement on systemic failures even as models improve.
This finding is not unique to Amazon’s improvement on the waimai dataset: in the 11 datasets we study, we observe the same pattern of homogeneous improvements from 2020-2022. In 4(b), we compare the observed improvement distribution [^4] The supplement § B.4 contains analysis comparing ‘net improvements’ to ‘potential improvements’ as well; the trends are consistent across both analyses. ($y$ axis) to the potential improvement distributions ($x$ axis) across all model improvements. We find a clear pattern: systemic failures (the [X, X] category) are represented less often in the observed improvement set than in the potential improvement set. This finding indicates that when models improve, they under-improve on users that are already being failed by other models. Instead, model improvements especially concentrate on instances where both other models succeeded already.
Ecosystem-level analysis in the context of model improvements disambiguates two plausible situations that are otherwise conflated: does single-model improvement (i) marginally or (ii) substantively reduce systemic failures? We find the reduction of systemic failures is consistently marginal: in every case, the reduction fails to match the the distribution seen in the previous year (i.e. every [X, X] red point is below the line in 4(b)).
@@ 5 Ecosystem-level Analysis in Dermatology (DDI)
Having demonstrated that ecosystem-level analysis reveals homogeneous outcomes across machine learning deployments, we apply the ecosystem-level methodology to medical imaging. We consider this setting to be an important use of ecosystem-level analysis because machine learning makes predictions that inform the high-stakes treatment decisions made by dermatologists.
@@ 5.1 Data
Daneshjou et al. (2022) introduced the Diverse Dermatology Images (DDI) dataset of 656 skin lesion images to evaluate binary classification performance at detecting whether a lesion was malignant or benign. Images were labelled with the ground-truth based on confirmation from an external medical procedure (biopsy). In addition, each image is annotated with skintone metadata using the Fitzpatrick scale according to one of three categories: Fitzpatrick I & II (light skin), Fitzpatrick III & IV (medium skin), and Fitzpatrick V & VI (dark skin). We use the predictions from Daneshjou et al. (2022) on DDI for two prominent ML models (ModelDerm (Han et al., 2020) and DeepDerm (Esteva et al., 2017)) and two board-certified dermatologists. [^5] Daneshjou et al. (2022) also evaluated a third model (HAM10K; Tschandl et al., 2018) that almost always predicts the majority class in this class-imbalanced setting. We exclude this model since its failures are not interesting, but replicate our analyses in the supplement Appendix C to show the findings still hold if it is included. We defer further details to Appendix C.
@@ 5.2 Results
Figure 5: Homogeneous outcomes for models and humans. Consistent with HAPI, model predictions (left) yield homogenous outcomes on DDI. Human predictions (right) are even more homogenous than models.
Finding 3: Both humans and models yield homogeneous outcomes; humans are more homogeneous.
We compare observed and baseline ecosystem-level outcomes on DDI for models (5(a)) and humans (5(b)). Consistent with the general trends in HAPI, model predictions yield homogeneous outcomes. For human predictions from board-certified dermatologists, we also see that outcomes are homogeneous. However, in comparing the two, we find that humans yield even more homogeneous outcomes. We take this to be an important reminder: while we predict that models are likely to produce homogeneous outcomes, we should also expect humans to produce homogeneous outcomes, and in some cases more homogeneous outcomes.
@@
Figure 6: Racial disparities for models but not humans. We stratify ecosystem-level analysis in DDI by the three skin tone categories, plotting the difference between observed and baseline rates. Models (left) show clear racial disparities, exhibiting the most homogeneity for the darkest skin tones, whereas humans (right) show no significant racial disparity.
@@
Standard analyses of machine learning’s impact focus on model performance across groups (e.g. Buolamwini and Gebru, 2018; Koenecke et al., 2020). In AI for medicine, several works have taken such an approach towards understanding fairness (Obermeyer et al., 2019; Seyyed-Kalantari et al., 2021; Kim et al., 2022; Colwell et al., 2022). Daneshjou et al. (2022) demonstrate racial disparities for both model and human predictions on DDI and find that predictive performance is worst for darkest skin tones, aligning with broader trends of racial discrimination in medicine and healthcare (Vyas et al., 2020; Williams and Wyatt, 2015; Fiscella and Sanders, 2016).
Ecosystem-level analysis can build on these efforts. We conduct the same analysis from Figure 5 stratified by skin tone. In previous experiments, we had observed systemic failures across the whole population. Here we measure systemic failures for subpopulations grouped by skin tone. In Figure 6, we plot the difference between the observed and baseline rates on the $y$ axis: the All bars (blue) reproduce the homogeneity results from Figure 5.
Strikingly, ecosystem-level-analysis surfaces an important contrast between model behavior (left) and human behavior (right). Models (6(a)) are most homogenous for darkest skin tones (Fitzpatrick V & VI; dark brown) and least homogeneous for the lightest skin tones (Fitzpatrick I & II; cream): The observed systemic failure rate for the darkest skin tones is 8.2% higher than the baseline, while for the lightest skin tones it is 1.5% lower than the baseline. By contrast, humans (6(b)) show no significant variation as a function of skin tone.
Ecosystem-level analysis therefore identifies a new form of racial disparity not previously documented. Critically, while prior works show racial disparities for both models and humans, here we find a form of racial disparity that is salient for models but not present for humans. We note that, in absolute terms, homogeneity is higher for all racial groups in human predictions than model predictions, though human predictions don’t display significant differences in homogeneity across racial group. This demonstrates that ecosystem-level analysis can reveal new dimensions of fairness, allowing stakeholders to identify the metrics that are most relevant in their context, be that model error or systemic failure rate. Our tool helps researchers and stakeholders evaluate the holistic ecosystem of algorithmic – and human – judgements that ultimately shapes outcomes for those subject to algorithmic judgement.
@@ 6.1 Explanations for Homogeneous Outcomes
We posit that “example difficulty” may give rise to homogeneous outcomes and provide three analyses that instantiate variants of this hypothesis.
@@
First, human ambiguity on the ground-truth may predict homogeneous outcomes. To test this, we make use of the ten human annotations per example in the fer+ dataset within HAPI. We find that the systemic failure rate is monotonically correlated with annotator disagreement with the majority label. This suggests that some systemic failures are correlated with the ambiguity or “hardness” of the image. However, we find that even some of the least ambiguous images, namely images on which all or most annotators agree, have systemic failures. . This indicates that human ambiguity is only partially explanatory of ecosystem-level model outcomes. We explore this further in § A.1.
@@
Second, human error may predict homogeneous outcomes. To test this, we compare human predictions on DDI with the ground truth biopsy results. We stratify ecosystem-level analysis by dermatologist performance, comparing (i) instances both dermatologists get right, (ii) instances they both get wrong, and (iii) instances where they disagree and exactly one is right. We find that when both dermatologists fail, there continues to be outcome homogenization. However, when both dermatologists succeed, there is no homogeneity and the observed rates almost exactly match the baseline rates for every image. This suggests human error is also partially predictive of ecosystem-level model outcomes. We explore this further in § A.2.
@@
Finally, more expressive theoretical models can better capture the observed trends than our simple full instance-level independence model. We introduce a two-parameter model. $\alpha$ fraction of instances are categorized as difficult and the remaining $1-\alpha$ are easy. A model’s failure rate $\bar{f_{j}}$ over all examples scales to $(1+\Delta)\bar{f_{j}}$ on hard examples and $\left(1-\frac{\alpha\Delta}{1-\alpha}\right)\bar{f_{j}}$ on easy examples. Partitioning examples in this way, while continuing to assume instance-level independence, inherently homogenizes: models are more likely to systemically succeed on easy instances and systemically fail on hard instances. To best fit the HAPI data, the resulting average $\alpha$ ($0.2-0.3$) and $\Delta$ ($1-4$, meaning these examples are 25×2-5$\times$2 - 5 × harder) values are quite extreme. In other words, for this theoretical model to fit the data, a significant fraction ($\approx 25\%$) would need to be considerably harder (3.5$\approx$ 3.5$\times$ 3.5 ×) than the overall performance. We explore this further in § A.3.
@@
These analyses contribute to an overarching hypothesis that example difficulty partially explains homogeneous outcomes. While we discuss the construct of difficulty in the supplement, we draw attention to two points. First, difficulty is largely in the eye of the beholder: what humans or models perceive as difficult can differ (e.g. adversarial examples; Goodfellow et al., 2014). Thus while example difficulty could be caused by inherent properties of the example (such as noisy speech or corrupted images), it could just as well be due to model properties, such as all the models having made similar architectural assumptions or having parallel limitations in their training data. Second, whether or not homogeneous outcomes are caused by example difficulty does not change their societal impact. The consequences of systemic failure can be material and serious (e.g. unemployment).
@@
Our paper shows that ecosystem-level research can reveal previously-invisible social impacts of machine learning. We believe this methodology concretizes the impact of decision-making in real contexts in which individuals would typically be affected by decisions from many actors (e.g. job applications, medical treatment, loan applications, or rent pricing). As we demonstrate in DDI, our methodology applies equally to human-only, machine-only, and more complex intertwined decision-making. We anticipate understanding of homogeneous outcomes arising from any or all of these sources will be valuable.
@@
Given the pervasive and persisting homogeneous outcomes we document, there may be a need for policy intervention. In many cases no single decision-maker can observe the decisions made by others in the ecosystem, so individual decision-makers (such as companies) may not know that systemic failures exist. In addition, systemic failures are not currently the responsibility of any single decision-maker, so no decision-maker is incentivized to act alone. Consequently, policy could implement mechanisms to better monitor ecosystem-level outcomes and incentivize ecosystem-level improvement. In parallel, regulators should establish mechanisms for recourse or redress for those currently systemically failed by machine learning (see Voigt and von dem Bussche, 2017; Cen and Raghavan, 2023).
@@ 7 Conclusion
We introduce ecosystem-level analysis as a new methodology for understanding the cumulative societal impact of machine learning on individuals. Our analysis on HAPI establishes general trends towards homogeneous outcomes that are largely unaddressed even when models improve. Our analysis on DDI exposes new forms of racial disparity in medical imaging that arise in model predictions but not in human predictions. Moving forward, we hope that future research will build on these empirical findings by providing greater theoretical backing, deeper causal explanations, and satisfactory sociotechnical mitigations. To ensure machine learning advances the public interest, we should use approaches like ecosystem-level analysis to holistically characterize its impact.
@@ Appendix A Data-centric Explanations for Homogeneous Outcomes
We test three variants of this hypothesis. In § A.1, we examine how the level of annotator disagreement in the ground truth label impacts ecosystem-level behavior. In § A.2, we test how the accuracy of human dermatologists in predicting the malignancy of a skin lesion image correlates with homogeneous outcomes. Finally, to build on these finer-grained empirical analyses, in § A.3, we introduce a more express theoretical model. Under this model, parameterized by two difficulty parameters, we compute a different baseline rate for ecosystem-level outcomes, showing it can better recover the observation distribution.